# Generative Modeling with Explicit Memory

### v1

@@ C.2 Hyperprarameters
We implement our model based on the original REPA implementation [70]. We use the AdamW optimizer [33] with a constant learning rate of $1\times 10^{-4}$, $\beta_{1}=0.9$, and $\beta_{2}=0.999$, without weight decay. To accelerate training, we employ mixed-precision (fp16) computation along with gradient clipping.
For latent space generation, we pre-compute compressed latent vectors from raw images using the Stable Diffusion VAE [55] and utilize these latent vectors as input. In contrast, for pixel space generation, we directly use the raw pixel data as input. Although we experimented with data augmentations such as flipping, we found that they did not significantly improve performance. Therefore, we do not apply any data augmentation in our experiments.
For projecting memory snippets into the backbone hidden dimension, we utilize a three-layer MLP with SiLU activations for diffusion transformers, following Yu et al. [70]. We provide detailed hyperparameter configurations in Tables 6 and 7.
@@ B L XL
Table 7: Training settings. We present the training settings for all models and training algorithms on the ImageNet 64$\times$64 dataset (left) and the ImageNet 256$\times$256 dataset (right).
@@
Batch size 128 128 128
Training iterations 400K 400k 400k
@@
Sampler Euler Euler Euler
Sampling steps 50 50 50
Classifier-free Guidance $\times$ $\times$ $\times$
@@
Input dim. 64$\times$64$\times$3 64$\times$64$\times$3 64$\times$64$\times$3
@@
Bank similarity measure Euclidean Euclidean Euclidean
Bank objective MSE MSE MSE
@@
Batch size 256 256 256
Training iterations 700K 700k 700k
@@
Sampler Euler Euler Euler
Sampling steps 50 50 50
Classifier-free Guidance $\times$ $\times$ $\times$
@@
Bank size 50k 50k 50k
Bank similarity measure Euclidean Euclidean Euclidean
Bank objective MSE MSE MSE
@@
We use Dinov2-B [48] as the encoder across all experiments, as it has been shown to significantly enhance the learning of better representations in diffusion models [70]. Dinov2-B offers superior performance, making it an ideal choice for facilitating the efficient training for constructing the memory bank.
@@
For the memory bank size, we set it to 50K for CIFAR-10 and 1.2M for ImageNet 64$\times$64 and ImageNet 256$\times$256. Note it is essential to maintain a relatively large memory bank for optimal performance, as demonstrated in Section A. Further results on the impact of memory bank size on performance are provided in Section A, where we detailed how varying the memory bank size affects the model’s efficacy.
@@
All models are primarily trained on NVIDIA H800 8-GPU setups, each equipped with 80GB of memory. The training speed for GMem-XL is approximately 1.15 seconds per 1KNI.
Appendix D Interpolation
D.1 GMem is Not a Memory Machine
@@
D.2 Interpolation Between Memory Snippets
To investigate this ability, we conduct an interpolation experiment. We use the ImageNet $256\times 256$ dataset and employ the model checkpoint obtained after 180M (NI), as described in  Table 7.
@@ D.3 Interpolation results
Appendix E Transferability of the Memory Bank
@@
E.1 Experimental Setup
To investigate the transferability, we trained a Memory Bank $\mathbf{M}_{\mathrm{CIFAR}}$ on the CIFAR-10 dataset and directly transferred it to a model trained on ImageNet 256 $\times$ 256 ${\textsc{GMem}}_{\mathrm{IN256}}$ to guide image generation. We used the checkpoint from the ImageNet model at 180M (NI) for generation. The detailed experimental settings are provided in  Table 6 and  Table 7.
E.2 Results
@@
Appendix F Text-to-Image Generation
@@
F.1 Experimental Setup
In our text-to-image experiments, we utilize BERT-base [11] as the text encoder, using class names as textual supervision signals and the CLS token as the textual feature representation. We employ the model trained on CIFAR-10 with 22.5M (NI) as the generator.
For mapping text features to the Memory Bank, we use a simple MLP with dimensions 768$\times$768. We train this mapping using the Adam optimizer with a learning rate of $1\times 10^{-4}$, optimizing the CLIP Loss [54] over 60 epochs.
F.2 Experimental Results
@@ Appendix G Derivation
G.1 Denoising Diffusion Probabilistic Models
Diffusion models [24] aim to model a target distribution $p(\mathbf{x})$ by learning a gradual denoising process that transitions from a Gaussian distribution $\mathcal{N}(0,\mathbf{I})$ to $p(\mathbf{x})$. The core idea is to learn the reverse process $p(\mathbf{x}_{t-1}|\mathbf{x}_{t})$ of a predefined forward process $q(\mathbf{x}_{t}|\mathbf{x}_{0})$, which incrementally adds Gaussian noise to the data starting from $\mathbf{x}_{0}\sim p(\mathbf{x})$ over $T$ time steps.
@@ G.1 Denoising Diffusion Probabilistic Models
In the DDPM framework introduced by Ho et al. [24], the reverse process $p(\mathbf{x}_{t-1}|\mathbf{x}_{t})$ is parameterized as:
@@
For the variance $\Sigma_{\theta}(\mathbf{x}_{t},t)$, Ho et al. [24] initially set it to $\sigma_{t}^{2}I$ with $\beta_{t}=\sigma_{t}^{2}$. However, Nichol and Dhariwal [47] demonstrated that performance improves when $\Sigma_{\theta}(\mathbf{x}_{t},t)$ is learned jointly with $\varepsilon_{\theta}(\mathbf{x}_{t},t)$. They propose optimizing the variational lower bound (VLB) objective:
@@
G.2 Stochastic interpolating
In contrast to DDPM, flow-based models [15, 41] address continuous time-dependent processes involving data samples $\mathbf{x}^{*}\sim p(\mathbf{x})$ and Gaussian noise $\varepsilon\sim\mathcal{N}(0,\mathbf{I})$ over the interval $t\in[0,1]$. The process is formulated as:
@@ G.2 Stochastic interpolating
According to Albergo et al. [1], stochastic interpolants satisfy the following conditions when $\alpha_{t}$ and $\sigma_{t}$ are chosen such that: 1. $\alpha_{t}^{2}+\sigma_{t}^{2}>0$ for all $t\in[0,1]$, 2. Both $\alpha_{t}$ and $\sigma_{t}$ are differentiable over the interval $[0,1]$, 3. Boundary conditions are met: $\alpha_{1}=\sigma_{0}=0$ and $\alpha_{0}=\sigma_{1}=1$.
@@
It’s noteworthy that existing score-based diffusion models, including DDPM [24], can be interpreted within an SDE framework. Specifically, their forward diffusion processes can be viewed as predefined (discretized) forward SDEs that converge to an equilibrium distribution $\mathcal{N}(0,\mathbf{I})$ as $t\to\infty$. Training is conducted over $[0,T]$ with a sufficiently large $T$ (e.g., $T=1000$) to ensure that $p(\mathbf{x}_{T})$ approximates an isotropic Gaussian. Generation involves solving the corresponding reverse SDE, starting from random Gaussian noise $\mathbf{x}_{T}\sim\mathcal{N}(0,\mathbf{I})$. In this context, $\alpha_{t}$, $\sigma_{t}$, and the diffusion coefficient $w_{t}$ are implicitly defined by the forward diffusion process, potentially leading to a complex design space in score-based diffusion models [28].

### v2

@@ B.2 Hyperprarameters
We implement our model based on the original REPA implementation (Yu et al., 2024). We use the AdamW optimizer (Kingma & Ba, 2014) with a constant learning rate of $1\times 10^{-4}$, $\beta_{1}=0.9$, and $\beta_{2}=0.999$, without weight decay. To accelerate training, we employ mixed-precision (fp16) computation along with gradient clipping.
For latent space generation, we pre-compute compressed latent vectors from raw images using the Stable Diffusion VAE (Rombach et al., 2022) and utilize these latent vectors as input. In contrast, for pixel space generation, we directly use the raw pixel data as input. Although we experimented with data augmentations such as flipping, we found that they did not significantly improve performance. Therefore, we do not apply any data augmentation in our experiments.
For projecting memory snippets into the backbone hidden dimension, we utilize a three-layer MLP with SiLU activations for diffusion transformers, following Yu et al. (2024). We provide detailed hyperparameter configurations in Tables 6 and 7.
@@ B L XL
Table 7: Training settings. We present the training settings for all models and training algorithms on the ImageNet 256$\times$256 dataset (left) and the ImageNet 512$\times$512 dataset (right).
@@
Batch size 1024 1024 1024
Training iterations 200K 200K 200K
@@
Sampler Heun Heun Heun
Sampling steps 100 100 100
@@
Input dim. 32$\times$32$\times$3 32$\times$32$\times$3 64$\times$64$\times$3
@@
Batch size 1024 1024 1024
Training iterations 500K 500k 500k
@@
Sampler Heun Heun Heun
Sampling steps 100 100 100
@@
Bank size 1.28M 1.28M 1.28M
@@
We use Dinov2-B (Oquab et al., 2023) as the encoder across all experiments, as it has been shown to significantly enhance the learning of better representations in diffusion models (Yu et al., 2024). Dinov2-B offers superior performance, making it an ideal choice for facilitating the efficient training for constructing the memory bank.
@@
For the memory bank size, we set it to 50K for CIFAR-10 and 1.2M for ImageNet 256$\times$256 and ImageNet 512$\times$512. Note it is essential to maintain a relatively large memory bank for optimal performance, as demonstrated in Section 5.4.
@@
All models are primarily trained on NVIDIA H800 8-GPU setups, each equipped with 80GB of memory. The training speed for GMem-XL is approximately 2.71 seconds per step.
Appendix C Interpolation
C.1 GMem is Not a Memory Machine
@@
C.2 Interpolation Between Memory Snippets
To investigate this ability, we conduct an interpolation experiment. We use the ImageNet $256\times 256$ dataset and employ the model checkpoint obtained after 140 epochs, as described in  Table 7.
@@ C.3 Interpolation results
C.4 Additional Applications
Figure 9: Ablation study on the variation of FID-50K with respect to number of training images (NI). We analyze the impact of NI during the training of GMem using different backbones on CIFAR-10 dataset. Specifically, while GMem-XL initially shows a higher FID than GMem-L at the early stages of training, after $1.75$M NI, GMem-XL consistently outperforms both GMem-L and GMem-B in terms of FID and maintains this lower value until convergence, where the final FID reaches $1.22$.
Transferable memory bank across datasets.
Appendix D demonstrates the transferability and generalization of the memory bank used to guide GMem across different datasets. Specifically, we train ${GMem}_{\textrm{IN64}}$ and ${GMem}_{\textrm{CIFAR}}$ models on ImageNet $64\times 64$ and CIFAR-10, respectively, resulting in memory banks of $\mathbf{M}_{\textrm{IN64}}$ and $\mathbf{M}_{\textrm{CIFAR}}$. We then directly apply $\mathbf{M}_{\textrm{IN64}}$ to guide the sampling process of ${GMem}_{\textrm{CIFAR}}$. Our results show that ${GMem}_{\textrm{CIFAR}}$ is still able to generate information consistent with the knowledge provided by $\mathbf{M}_{\textrm{IN64}}$.
Application to T2I generation.
This integration is achieved through a two-step process:
  1. (a)
    Pretraining the mapping function: We pretrain a mapping function $\boldsymbol{\psi}$ that transforms text prompt distributions $\mathbf{p}({\text{text}})$ into memory snippet distributions $\mathbf{p}(\mathbf{s})$ using a simple contrastive loss (Radford et al., 2021).
  2. (b)
Comprehensive descriptions of the T2I generation process, experimental setup, and the resulting generated images are provided in Appendix D.
Appendix D Transferability of the Memory Bank
@@
D.1 Experimental Setup
To investigate the transferability, we trained a Memory Bank $\mathbf{M}_{\mathrm{CIFAR}}$ on the CIFAR-10 dataset and directly transferred it to a model trained on ImageNet 256 $\times$ 256 ${{GMem}}_{\mathrm{IN256}}$ to guide image generation. We used the checkpoint from the ImageNet model at 140 epochs for generation. The detailed experimental settings are provided in  Table 6 and  Table 7.
D.2 Results
@@
Appendix E Text-to-Image Generation
@@
E.1 Experimental Setup
In our text-to-image experiments, we utilize BERT-base (Devlin, 2018) as the text encoder, using class names as textual supervision signals and the CLS token as the textual feature representation. We employ the model trained on CIFAR-10 with 450 epochs as the generator.
For mapping text features to the Memory Bank, we use a simple MLP with dimensions 768$\times$768. We train this mapping using the Adam optimizer with a learning rate of $1\times 10^{-4}$, optimizing the CLIP Loss (Radford et al., 2021) over 60 epochs.
E.2 Experimental Results
@@ Appendix F Derivation
F.1 Denoising Diffusion Probabilistic Models
Diffusion models (Ho et al., 2020) aim to model a target distribution $p(\mathbf{x})$ by learning a gradual denoising process that transitions from a Gaussian distribution $\mathcal{N}(0,\mathbf{I})$ to $p(\mathbf{x})$. The core idea is to learn the reverse process $p(\mathbf{x}_{t-1}|\mathbf{x}_{t})$ of a predefined forward process $q(\mathbf{x}_{t}|\mathbf{x}_{0})$, which incrementally adds Gaussian noise to the data starting from $\mathbf{x}_{0}\sim p(\mathbf{x})$ over $T$ time steps.
@@ F.1 Denoising Diffusion Probabilistic Models
In the DDPM framework introduced by Ho et al. (2020), the reverse process $p(\mathbf{x}_{t-1}|\mathbf{x}_{t})$ is parameterized as:
@@
For the variance $\Sigma_{\theta}(\mathbf{x}_{t},t)$, Ho et al. (2020) initially set it to $\sigma_{t}^{2}I$ with $\beta_{t}=\sigma_{t}^{2}$. However, Nichol & Dhariwal (2021) demonstrated that performance improves when $\Sigma_{\theta}(\mathbf{x}_{t},t)$ is learned jointly with $\varepsilon_{\theta}(\mathbf{x}_{t},t)$. They propose optimizing the variational lower bound (VLB) objective:
@@
F.2 Stochastic interpolating
In contrast to DDPM, flow-based models (Esser et al., 2024; Liu et al., 2023) address continuous time-dependent processes involving data samples $\mathbf{x}^{*}\sim p(\mathbf{x})$ and Gaussian noise $\varepsilon\sim\mathcal{N}(0,\mathbf{I})$ over the interval $t\in[0,1]$. The process is formulated as:
@@ F.2 Stochastic interpolating
According to Albergo et al. (2023), stochastic interpolants satisfy the following conditions when $\alpha_{t}$ and $\sigma_{t}$ are chosen such that: 1. $\alpha_{t}^{2}+\sigma_{t}^{2}>0$ for all $t\in[0,1]$, 2. Both $\alpha_{t}$ and $\sigma_{t}$ are differentiable over the interval $[0,1]$, 3. Boundary conditions are met: $\alpha_{1}=\sigma_{0}=0$ and $\alpha_{0}=\sigma_{1}=1$.
@@
It’s noteworthy that existing score-based diffusion models, including DDPM (Ho et al., 2020), can be interpreted within an SDE framework. Specifically, their forward diffusion processes can be viewed as predefined (discretized) forward SDEs that converge to an equilibrium distribution $\mathcal{N}(0,\mathbf{I})$ as $t\to\infty$. Training is conducted over $[0,T]$ with a sufficiently large $T$ (e.g., $T=1000$) to ensure that $p(\mathbf{x}_{T})$ approximates an isotropic Gaussian. Generation involves solving the corresponding reverse SDE, starting from random Gaussian noise $\mathbf{x}_{T}\sim\mathcal{N}(0,\mathbf{I})$. In this context, $\alpha_{t}$, $\sigma_{t}$, and the diffusion coefficient $w_{t}$ are implicitly defined by the forward diffusion process, potentially leading to a complex design space in score-based diffusion models (Karras et al., 2022).